# Breaking Language Barriers or Reinforcing Bias? A Study of Gender and Racial Disparities in Multilingual Contrastive Vision–Language Models

### v1

@@ 4 Results and Analysis
Table 2: Maximum skew values for gender and race stereotypes across ten languages. For each model–dataset pair (mCLIP, NLLB-CLIP, CAPIVARA-CLIP on FairFace and PATA), we report the maximum skew for crime ($\max s_{G}^{c},\max s_{R}^{c}$), communion ($\max s_{G}^{com},\max s_{R}^{com}$), and agency ($\max s_{G}^{ag},\max s_{R}^{ag}$) submetrics.
Low-resource languages
As shown in Figure 1 and detailed in Tables 2 and 3, bias intensifies where captions are scarcest. Across Hindi, Xhosa and Portuguese, CAPIVARA registers the largest average gender–crime skew ($1.66$); mCLIP is lowest at $0.29$. Within languages the picture fragments. In Hindi, NLLB reaches a race–crime skew of $3.86$, dwarfing mCLIP ($2.71$) and CAPIVARA ($2.42$). Xhosa shows the highest race–agency skew under CAPIVARA ($0.84$), while Portuguese shifts the maximum race–crime value to NLLB ($3.10$) and the peak gender–agency value to CAPIVARA ($0.74$).
(a) Gender-related skews
(b) Race-related skews
Figure 1: Bias skew metrics for low-resource languages (Hindi, Xhosa, Portuguese). The left panel shows gender-related skews and the right panel shows race-related skews for mCLIP, NLLB-CLIP, and CAPIVARA-CLIP. Scarce caption data amplify crime and agency stereotypes most strongly in CAPIVARA-CLIP, while mCLIP remains comparatively lower in overall skew.
High-resource languages
With abundant captions the ranking changes. On English, French and Spanish images the race–crime metric is governed by mCLIP ($3.05$ on average), whereas CAPIVARA drops to $2.45$. The worst gender–agency scores, however, move from model to model: they occur under NLLB in English ($0.20$), shift to CAPIVARA in French ($0.44$) and remain with CAPIVARA in Spanish ($0.47$)(Figure 2 alongside Tables 2 and 3).
(a) Gender metrics (balance, agency, crime, average).
(b) Race metrics (mean, agency, crime, average).
Figure 2: Bias skew metrics in high-resource languages (English, French, Spanish) across multilingual CLIP variants. Though caption data are ample, all models accentuate crime-related skews relative to monolingual baselines, and agency/community biases vary by architecture.
@@
Grammatical gender amplifies stereotypes. In the gender-neutral trio (Turkish, Farsi, Finnish) the average gender–crime skew is almost negligible for mCLIP ($0.06$) but rises to $3.11$ for NLLB, illustrating how a shared encoder can import English biases (Figure 4. When the language itself is highly gendered (Spanish, French, Slovak) all models deteriorate: the mean gender–crime skew reaches $2.47$ for mCLIP, $2.82$ for NLLB and $2.21$ for CAPIVARA. Spanish is the most extreme case, with CAPIVARA hitting $3.32$ gender–crime and $4.11$ race–communion skew (Figure 3, and Tables 2, 3).
(a) Gender-related skews (balance, agency, crime, average).
(b) Race-related skews (mean race, agency, crime, average).
Figure 3: Bias skew metrics for highly gendered languages (Spanish, French, Slovak) across multilingual CLIP variants. The left panel shows gender-related skews and the right panel shows race-related skews, illustrating how grammatical gender amplifies different stereotype axes.
(a) Gender-skew components for each multilingual CLIP variant in Turkish, Farsi, Finnish.
(b) Race-skew components for each multilingual CLIP variant in Turkish, Farsi, Finnish.
Figure 4: Bias skew metrics for gender-neutral languages (Turkish, Farsi, Finnish). The left panel shows gender-skew components, and the right panel shows race-skew components, illustrating how a shared encoder (NLLB-CLIP) can import English-centric biases into languages without grammatical gender marking, whereas adapter-based models (mCLIP, CAPIVARA-CLIP) largely suppress these skews.
@@
Crime associations are exceptionally persistent: every multilingual model surpasses its own English baseline on crime skew in every language. Agency and communion, in contrast, act like architectural fingerprints. CAPIVARA shows the largest agency skew in low-resource languages, whereas mCLIP dominates communion skew in highly gendered contexts, reaching $2.97$ in French.
@@
Better retrieval often comes with worse fairness. CAPIVARA gains +3.4 R@1 in Portuguese relative to the OpenCLIP B/32 baseline, yet its gender–crime skew rises from $0.22$ to $1.05$. A similar pattern appears in Spanish: mCLIP improves R@1 by 4.2 points but lifts race–crime skew from $0.26$ to $3.39$. Low-resource recall, in short, is currently purchased at a substantial cost to equity.
Parameter-efficient multilingual CLIP models strengthen, rather than weaken, social stereotypes, and the sharpest increases appear precisely in the languages and bias axes where benchmarks are still rare. Fine-grained, language-aware reporting is therefore essential for any claim of fairness.
@@ 5 Discussion
Our audit shows that extending CLIP to additional languages increases retrieval accuracy but just as reliably enlarges the model’s capacity to reproduce social stereotypes. Gender and agency biases grow even in English, and the rise is steeper in every other language we examined. The pattern is not random: the largest surges occur precisely where textual resources are scarcest and where grammatical gender pervades the lexicon. Hindi and Xhosa, represented by the thinnest caption corpora, experience the most pronounced race and gender skew, while Spanish and French, whose morphology forces gender marking, amplify all stereotype categories.
Architecture further modulates these effects. A single shared encoder, as adopted by NLLB-CLIP, transports English gender stereotypes wholesale into gender-neutral languages. Designs that keep language-specific adapters—for instance CAPIVARA-CLIP—largely shield those languages, although their protection comes at the cost of greater bias in the truly low-resource tongues they rely on most. Morphological cues thus act as amplifiers or dampers depending on whether cross-lingual weight sharing re-imports external bias.
Not all stereotypes behave alike. Criminality associations are persistent across models and languages, suggesting deep-rooted artefacts in web imagery and captions. By contrast, agency and communion biases vary with architectural choices, indicating that regularisation or debiasing objectives could target them specifically. Because such divergences are obscured when scores are averaged over languages or stereotype axes, global leaderboards can mislead practitioners: a checkpoint that looks benign in English may still be sharply prejudiced for Hindi, Xhosa or any other under-audited language.
The practical implication is straightforward: fairness cannot be assumed from accuracy. Safe deployment of multilingual vision-language systems demands bias monitoring and mitigation during training—through balanced data curation, fairness-aware loss terms or counterfactual augmentation—rather than relegating those concerns to post-hoc fixes.
@@ 6 Conclusion
This paper delivered the first systematic, multi-axis audit of parameter-efficient multilingual CLIP variants, spanning languages that differ sharply in data availability and grammatical gender. Our analysis shows that scaling CLIP to new languages does not inoculate the model against social stereotypes; on the contrary, multilingual adaptations consistently intensify them. Biases grow in English and grow faster elsewhere, with the steepest increases observed in low-resource and morphologically gendered languages. Architectural choices also matter: a single shared encoder, as in NLLB-CLIP, transfers English gender stereotypes wholesale into languages whose grammar lacks gender marking, whereas more loosely coupled designs leave those same languages comparatively untainted. The interaction between language morphology and encoder sharing therefore determines the amplitude of the resulting bias.
@@
7 Limitations
This study is constrained in several respects that should temper the generality of its claims. First, both FairFace and the PATA stereotype suite embed North-American notions of race and personhood; consequently, the bias patterns we report may not align with local taxonomies elsewhere. Second, our principal metric, max-skew, foregrounds extreme probability gaps but says little about subtler distributional shifts or downstream error types such as false negatives. Third, all prompt templates were translated with GPT o3 and back-validated; any translation errors or biases in that model propagate into our measurements, with heightened impact on lower-resource languages. Fourth, although the ten languages analysed cover four families, they omit right-to-left scripts and ultra-low-resource languages for which machine-translation noise is substantially higher. Fifth, we evaluate only five public checkpoints under a zero-shot protocol; larger proprietary backbones, vision–language pipelines with re-ranking, or models fine-tuned with fairness objectives could behave differently. Finally, all conclusions inherit the biases of the underlying datasets and the selection of stereotype axes, which do not yet include attributes such as age, disability, or caste. These caveats underscore the need for continued, locale-aware auditing as multilingual vision–language technology evolves.
Appendix A Corpus-level harm rates

### v3

@@ 4 Results and Analysis
SIGLIP2 $F$ $\max s_{R}^{c}$ 1.22 0.59 4.40 0.21 5.58 5.40 0.55 1.93 1.76 42.12
$\max s_{R}^{com}$ 0.26 0.40 0.42 0.15 0.06 0.55 0.23 0.23 0.68 0.70
$\max s_{R}^{ag}$ 0.22 0.11 0.23 0.10 0.44 0.37 1.02 0.04 0.55 0.14
PT $\max s_{R}^{c}$ 10.66 6.17 12.05 0.47 2.01 30.02 6.12 3.96 1.20 1.74
$\max s_{R}^{com}$ 0.57 0.37 0.14 0.12 0.42 0.18 0.66 0.25 0.36 0.16
$\max s_{R}^{ag}$ 0.17 0.33 0.58 0.36 0.79 0.32 0.28 1.20 0.27 1.75
Table 4: Cross‑lingual race bias (mean pairwise max‑skew). Average of all unordered race‑pair skews ($\max s_{R}$) per axis across ten languages. This corpus‑level view mitigates single–class outliers and highlights systematic disparities. Higher numbers indicate stronger association gaps; coloured entries mark the worst axis for each model, red for $\max s_{R}^{c}$, orange for $\max s_{R}^{\text{com}}$, and green for $\max s_{R}^{\text{ag}}$.
Low‑resource languages
Figures 1,2 and Tables 24 bias intensifies where captions are scarcest. Across Hindi, Xhosa and Portuguese the average gender–crime skew is highest for SigLIP2 ($1.53$) followed by CAPIVARA ($0.61$) and lowest for NLLB‑CLIP ($0.11$), with mCLIP ($0.26$) in between. Within languages the leaders shift: in Hindi race–crime peaks under SigLIP2 ($5.40$), Xhosa gives the race–agency maximum to CAPIVARA ($2.31$), and Portuguese keeps race–crime highest for NLLB ($0.75$). Symmetric KL tracks these skews—broadest for CAPIVARA in Xhosa ($\mathrm{SKL}_{c}=0.15$) and near‑zero for mCLIP. Corpus‑level harm rates mirror the gradient: harmful top‑1 predictions (%NC + %C) exceed 80 % for SigLIP2 and NLLB in Xhosa, but stay below 30 % for mCLIP, underscoring how bias magnifies when both data and debiasing signals are sparse.
(a) Low‑resource languages
(b) High‑resource languages
Figure 1: Gender max‑skew on FairFace. (a) Low‑resource languages (hi, xh, pt). (b) High‑resource languages (en, es, fr). Bars show crime, communion and agency skews for the four multilingual checkpoints. Spikes for CAPIVARA in Xhosa and SigLIP2 in Hindi reveal how data scarcity can inflate gender–crime associations even when corresponding English skews remain modest.
High‑resource languages
With abundant captions the ranking changes. Across English, French and Spanish, the race–crime skew is now led by SigLIP2 (avg. $4.37$), trailed by NLLB‑CLIP ($2.58$), while mCLIP ($0.98$) and CAPIVARA ($0.60$) stay lower. Gender–agency outliers still rotate: they appear under NLLB in English ($0.20$), shift to CAPIVARA in French ($\mathbf{1.06}$), and remain with CAPIVARA in Spanish ($0.47$) (Figures 1,2 and Tables 24). Symmetric KL confirms the pattern: CAPIVARA shows the widest gender–agency gap in French ($\mathrm{SKL}_{\text{ag}}=0.52$), whereas SigLIP2 stays near zero on every axis. Corpus‑level harm rates follow suit: harmful top‑1 predictions (%NC + %C) peak for NLLB in English (74 %) but fall below 10 % for CAPIVARA across all three languages, placing the latter as the least toxic despite its high skew on specific traits.
(a) Low‑resource languages
(b) High‑resource languages).
Figure 2: Race mean‑max‑skew on FairFace. (a) Low‑resource languages (hi, xh, pt). (b) High‑resource languages (en, es, fr). Mean‑max‑skew averages disparities over all race pairs; the tallest bars confirm that race–crime stereotypes intensify under the shared‑encoder (NLLB‑CLIP) in Hindi and explode for SigLIP2 in Xhosa.
@@
Grammatical gender amplifies stereotypes. In the gender-neutral trio (Turkish, Farsi, Finnish) the average gender–crime skew is almost negligible for mCLIP ($0.06$) but rises to $3.11$ for NLLB and $2.38$ for SigLIP2, illustrating how a shared encoder can import English biases (Figure 3). When the language itself is highly gendered (Spanish, French, Slovak) all models deteriorate: the mean gender–crime skew reaches $2.47$ for mCLIP, $2.82$ for NLLB and $2.21$ for CAPIVARA. Spanish is the most extreme case, with CAPIVARA hitting $3.32$ gender–crime and $4.11$ race–communion skew (Figures 3,4 and Tables 24).
(a) Gender‑neutral languages
(b) Highly gendered languages
Figure 3: Gender max‑skew on FairFace by grammatical system. (a) Gender‑neutral languages (tr, fa, fi). (b) Highly gendered languages (es, fr, sl). Replacing CLIP’s text tower with a shared multilingual encoder (NLLB‑CLIP) leaves gender‑neutral skews small, whereas adapter‑based CAPIVARA and Web‑scale SigLIP2 show sharp increases once overt grammatical gender is present.
(a) Gender‑neutral languages
(b) Highly gendered languages
Figure 4: Race mean‑max‑skew on FairFace by grammatical system. (a) Gender‑neutral languages (tr, fa, fi). (b) Highly gendered languages (es, fr, sl). Race skews rise most for the loosely coupled CAPIVARA adapters in gender‑neutral Turkish ($3.25$) and for SigLIP2 in gendered French ($5.58$), underscoring that grammatical gender can interact with race biases in non‑obvious ways.
@@
Crime remains the most stubborn axis: every multilingual variant exceeds its own English baseline in every language, and the worst offender often shifts with data scarcity—NLLB‑CLIP in Hindi ($5.40$), CAPIVARA in Xhosa ($1.77$), SigLIP2 in Farsi ($3.16$). Communion behaves like a proxy for model size and filtering: SigLIP2 posts the lowest skews almost everywhere (<0.40<\!0.40 for 28 / 30 language–dataset pairs), whereas mCLIP reaches the ceiling in gender‑rich French ($2.97$). Agency is an architectural fingerprint: CAPIVARA dominates low‑resource agency skew (e.g. $2.31$ in Xhosa), while NLLB owns the English spike ($2.58$); SigLIP2 stays close to zero on this axis but shows the largest KL divergence for crime in Turkish ($\mathrm{SKL}_{c}=0.41$), revealing a thinner yet sharper bias tail.
@@
The fairness–accuracy trade‑off is axis‑specific. Parameter‑efficient methods improve recall at a clear cost: CAPIVARA gains +3.4+3.4 R@1 on Portuguese yet raises gender–crime skew from $0.22$ to $1.05$; mCLIP adds +4.2+4.2 R@1 in Spanish while race–crime climbs from $0.26$ to $3.39$. Encoder replacement (NLLB) keeps English recall intact but triggers double‑digit race–crime skews in gender‑neutral Finnish ($17.27$). Full Web‑scale re‑training (SigLIP2) Pareto‑dominates the other variants on agency and communion—cutting average skews by 70 % with no loss in XTD‑11 R@1—but leaves crime bias largely unsolved and can even explode it when captions are scarce (e.g. Xhosa, $42.12$).
In sum, no single architecture is bias‑free: crime stereotypes persist, agency and communion reflect modelling choices, and the sharpest disparities still surface in the very languages where evaluation data are thin. Fine‑grained, language‑aware reporting therefore remains indispensable for any fairness claim.
@@ 5 Discussion
Our audit confirms a persistent pattern: extending CLIP to new languages raises retrieval accuracy but just as reliably enlarges the model’s capacity to reproduce social stereotypes. Gender and agency bias grow even in English and rise steeply elsewhere. The surges are largest where captions are scarce (Hindi, Xhosa) and where grammar forces gender marking (Spanish, French), implicating resource imbalance and morphology as bias magnifiers.
Debiasing in the wild.
Web‑scale retraining with bias filtering does help—in SigLIP 2, agency and communion skews fall by up to 70 % and their symmetric‐KL values approach zero. Yet crime stereotypes remain stubborn: SigLIP 2 still exceeds every other checkpoint on race–crime in the sparsest language (42.1 in Xhosa) and barely improves on English. Large‑scale debiasing thus attenuates the “shallow” traits it targets, but it cannot erase deeper associations that are already baked into the English‑centric alignment space.
The centre–periphery anchor.
All four models share a hub‑and‑spoke geometry in which an English‑dominated embedding space serves as the universal alignment target. Whether through distillation (M‑CLIP), LoRA tuning (CAPIVARA), encoder replacement (NLLB) or full retraining (SigLIP 2), every new language is forced to map onto that biased centre. When captions are noisy or synthetic the mapping error shows up as stereotype skew: the shared encoder of NLLB imports English gender bias into gender‑neutral Turkish and Finnish; adapter‑based CAPIVARA protects those languages but inflates bias in its low‑resource targets; SigLIP 2 reduces most traits yet inherits the anchor’s crime associations, especially under severe data sparsity.
Bias axes are unequal.
Crime associations are exceptionally persistent: each model surpasses its English baseline on crime skew in every language. Agency and communion, by contrast, act like architectural fingerprints—worst for CAPIVARA in low‑resource settings, for NLLB in English, and lowest for SigLIP 2 almost everywhere. Because these divergences vanish when metrics are averaged, global leaderboards can mislead: a checkpoint that looks benign in English may be sharply prejudiced for Hindi, Xhosa or any under‑audited language.
Fairness versus accuracy.
Parameter‑efficient variants raise recall at a clear cost: CAPIVARA gains +3.4+3.4 R@1 in Portuguese but lifts gender–crime from $0.22$ to $1.05$; M‑CLIP adds +4.2+4.2 R@1 in Spanish yet drives race–crime from $0.26$ to $3.39$. SigLIP 2 Pareto‑dominates on agency and communion but leaves crime bias largely unsolved, illustrating that fairness improvements can be axis‑specific and that debiasing alone cannot offset the centre–periphery effect.
Taken together, our results underscore four overarching lessons. First, multilingual coverage alone does not guarantee equitable behavior. Second, the English‑centric anchor that underlies current alignment spaces is a primary amplifier of bias, especially in low‑resource settings. Third, meaningful debiasing must reach beyond simple data filtering and reshape the alignment geometry itself—for instance, by introducing multiple language pivots or language‑specific anchors. Fourth, any credible claim of fairness in multilingual vision–language systems requires fine‑grained, language‑aware evaluation and reporting.
@@ 6 Conclusion
This paper delivered the first systematic, multi‑axis audit of parameter‑efficient multilingual CLIP variants, spanning languages that differ sharply in data availability and grammatical gender. Our analysis shows that scaling CLIP to new languages does not inoculate the model against social stereotypes; on the contrary, multilingual adaptations consistently intensify them. Biases grow in English and grow faster elsewhere, with the steepest increases observed in low‑resource and morphologically gendered languages. Architectural choices also matter: a single shared encoder, as in NLLB-CLIP and SIGLIP2, transfers English gender stereotypes wholesale into languages whose grammar lacks gender marking, whereas more loosely coupled designs leave those same languages comparatively untainted. The interaction between language morphology and encoder sharing therefore determines the amplitude of the resulting bias. Nevertheless, prevailing debiasing strategies provide only partial relief: although they curb agency‑ and communion‑related skews, they leave crime‑linked stereotypes largely intact—and in caption‑sparse settings such as Xhosa, they can even exacerbate them. Thus, underscoring the need for deeper, language‑aware mitigation approaches.
@@
Limitations
Our findings should be interpreted in light of several constraints.
Data provenance. Both FairFace and the PATA stereotype suite encode a North-American taxonomy of race and personhood; bias patterns may differ under region-specific classifications of caste, tribe or ethnicity. The datasets also lack annotations for disability, age and religion, so we cannot assess those axes. More broadly, vision–language datasets and racial categories are not culturally uniform, and our results should be read with this caveat in mind.
Cultural validity of templates and labels. Our audit queries whether, for the same face under zero-shot classification, a model assigns a higher score to a negative attribute label (e.g., “a photo of a criminal”) than to the correct demographic description, using a small candidate set comprising the demographic label, the negative attribute label, and non-human distractors (to ensure the model is not simply defaulting to any human label). We do not employ a bespoke cultural-adaptation protocol; templates were machine-translated and, where possible, validated by native speakers for fluency and neutrality (e.g., Hindi). We will release the full per-language label translations in the appendix and emphasize that our cross-lingual comparisons are made within a fixed, documented label set rather than across culturally contingent taxonomies.
Metric scope. We focus on max-skew and symmetric-KL, which highlight extreme probability gaps but shed little light on false-negative error modes or long-tail distributional shifts—important for safety-critical deployments.
Mechanistic understanding of bias propagation and transfer. Our emphasis is an empirically grounded, cross-lingual audit—covering four architectures, ten languages, and multiple bias axes with a fully reproducible toolkit—rather than a causal/mechanistic study. We therefore stop short of causal claims. As concrete next steps, we outline embedding- subspace analyses, causal/interventional tests, and controlled synthetic datasets—directions that our released audit harness is designed to enable.
Prompt translation. Templates were machine-translated with GPT o3 and then human-validated by bilingual speakers. Any residual inaccuracies or lexical bias may propagate into the results, especially for ultra-low-resource tongues.
Language coverage. Our ten languages span four families but exclude right-to-left scripts (Arabic, Hebrew), logographic writing (Chinese) and very low-resource languages where translation noise is higher and caption corpora are thinner.
Model selection and audited scope. We audit four widely used public checkpoints under a zero-shot protocol, including SigLIP-2—a debiased training recipe that in our scarce-caption settings still exhibits persistent crime associations. Larger proprietary backbones, re-ranked retrieval pipelines, or models fine-tuned with fairness-aware objectives could behave differently; likewise, we do not test vision encoders other than ViTs. We position such architectural exploration (e.g., decentralizing English-hub alignment, fairness-aware training objectives) as future work. To facilitate this, we release our evaluation toolkit and prompt inventories so new architectures can be plugged into the exact same audit protocol.
Architectural bias. All models share a centre–periphery design anchored in an English space. Our study cannot disentangle whether observed skews stem from that geometry or from dataset imbalance alone; exploring multi-pivot or language-specific anchors is left for future work.
These limitations underscore the need for ongoing, locale-aware auditing as multilingual vision–language systems evolve and as new datasets and mitigation strategies emerge.
Acknowledgements
MP was supported by the UK Engineering and Physical Sciences Research Council via Responsible AI UK (grant number EP/Y009800/1, project KP0016, “AdSoLve: Addressing Socio-technical Limitations of LLMs for Medical and Social Computing”); by the European Union’s Horizon Europe research and innovation programme under grant agreement number 101214398 (ELLIOT); and by the Slovenian Research and Innovation Agency (ARIS) through the Gravitacije project LLM4DH (“Large Language Models for Digital Humanities”, GC-0002), the project CroDeCo (“Cross-Lingual Analysis for Detection of Cognitive Impairment in Less-Resourced Languages”, J6-60109) and the research programme “Knowledge Technologies” (P2-0103). Views and opinions expressed are however those of the author(s) only and do not necessarily reflect those of the European Union or the European Commission. Neither the European Union nor the European Commission can be held responsible for them. ZA is supported by Google DeepMind PhD Fellowship and thanks their Google DeepMind mentor, David Stutz, for guidance and support.
We thank the bilingual volunteers for their careful review and validation of the LLM-generated translations: Seyedpeyman Hosseini (Farsi); Sarah Chamouni (French); Sadaf Saiyed (Hindi); Tiia Pelkonen and Christian Guckelsberger (Finnish); Mustafa Işık (Turkish); Onelisa Slater and Eva-Marie Bloom Ström (Xhosa); Rafael Frade (Portuguese, Spanish); and Teo Radetic (Slovenian).
Appendix A Methods Supplementary Material
This section supplies details model comparisons, data description, full formalism, and prompt inventories that were abridged in §3. Table, figure and equation numbers are local to the appendix.
A.1 Models Details
A CLIP‑style model consists of two towers: a vision encoder
f v   that maps an image
x to an embedding
v R d and a text encoder
f t   that sends a caption
t to
u R d . During inference an image and its paired caption should have high cosine similarity, whereas mismatched pairs should be far apart. We compare multilingual checkpoints that retain the original vision encoder—hence preserving zero‑shot object recognition—while adapting or re‑training the text encoder for many languages.
Evaluation metric.
To place all models on the same scale we report recall at ten (R@10) on the Crossmodal‑3600 benchmark (Thapliyal et al., 2022), unless the source paper publishes only R@1, in which case we state that explicitly. R@10 is the fraction of queries whose correct match is found among the ten highest‑scoring candidates.
M‑CLIP
(Carlsson et al., 2022). M‑CLIP is a post‑hoc multilingual extension of OpenAI CLIP. It freezes the ViT‑B/32 vision tower and replaces the English text encoder with an XLM‑R Base Base   network trained by teacher–student distillation. English captions from MS‑COCO, GCC and VizWiz are translated into 68 languages with Marian MT; the student minimises the mean‑squared error to the teacher’s embeddings, so no images are required at training time. Despite this text‑only optimisation, M‑CLIP reaches 79.8 R@10 on Crossmodal‑3600, essentially matching English CLIP while supporting dozens of new languages.
NLLB‑CLIP
(Visheratin, 2023b). NLLB‑CLIP swaps CLIP’s text tower for Meta’s 3.3 B‑parameter No Language Left Behind encoder, again keeping the ViT‑B/32 image tower frozen. Fine‑tuning uses 106 k LAION–COCO images whose captions are automatically translated into all 201 FLORES‑200 languages, thereby maximising linguistic breadth. The resulting model attains 81.2 R@10 (and 43.4 R@1) on Crossmodal‑3600, with particularly large gains in low‑resource languages such as Quechua and Māori.
CAPIVARA‑CLIP
(dos Santos et al., 2023). CAPIVARA targets rapid adaptation to a single low‑resource language. It first re‑generates English captions for CC3M images with BLIP‑2, translates them into Portuguese, Hindi and Xhosa, and then fine‑tunes only the text encoder of an OpenCLIP ViT‑B/32 checkpoint via lightweight LoRA adapters under the LiT objective. Two GPU‑hours of training lift Portuguese retrieval on MS‑COCO from 59 R@10 to 80.3 R@10 while leaving English accuracy intact, demonstrating an inexpensive path to domain‑specific multilinguality.
SigLIP 2
(Tschannen et al., 2025). SigLIP 2 is trained from scratch on WebLI, a 10 B‑image, 12 B‑caption corpus in 109 languages. Its curriculum interleaves the Sigmoid contrastive loss with decoder‑based captioning and masked‑token prediction, and applies the “Clip the Bias” filtering pipeline (Ibrahim Alabdulmohsin et al., 2024) to reduce demographic bias. The public ViT‑L/16 checkpoint reaches 84.7 R@10 on Crossmodal‑3600—currently the best open‑weights score—while cutting female representation bias in FairFace from 35
English baselines.
For reference we also audit the original OpenAI CLIP (ViT‑L/14, ViT‑B/32) and OpenCLIP (ViT‑B/32) checkpoints, which establish an upper bound on English retrieval (up to 88 R@10 on MS‑COCO) and a lower bound on multilingual fairness.
A.2 Language Partitions
low-resource $:\{\text{pt},\text{xh},\text{hi}\},$ high-resource $:\{\text{en},\text{es},\text{fr}\},$
gender-less $:\{\text{tr},\text{fa},\text{fi}\},$ gender-rich $:\{\text{sl},\text{es},\text{fr}\}.$
A.3 Datasets Details
FairFace.
FairFace is a large-scale face dataset containing 108 501 celebrity‐free portraits sampled from Flickr under a CC BY-NC licence (Karkkainen and Joo, 2021). Each image is annotated with binary gender (male/female) and one of seven self-identified race categories: {White, Black, Indian, East-Asian, South-East-Asian, Middle-Eastern, Latino} To obtain a balanced validation set, we draw $N_{r,g}\approx 782$ images per race–gender cell, yielding 7 races × 2 genders × 782   10 954 portraits.
Protected-Attribute Tag Association (PATA).
The PATA benchmark comprises 4 934 face images annotated for bias measurement in vision–language models (Seth et al., 2023). Each portrait carries binary gender labels (male/female) and one of five ethno-racial identities: {Black, Caucasian, East-Asian, Hispanic/Latino, Indian} We retain only gender and race annotations (dropping age) and evaluate on the official test split to ensure direct comparability with prior work.
A.4 Complete Metric Definitions
Pairwise Skew.
For any two groups $A,B\in G$ we quantify their relative disparity as
$\mathrm{maxs}(A,B)=\max\!\biggl(\frac{|p_{A}-p_{B}|}{p_{A}},\,\frac{|p_{A}-p_{B}|}{p_{B}}\biggr),$ (2)
with $p_{A}=s(A,c)$ and $p_{B}=s(B,c)$. Skew is scale‑free: if $A$ scores twice as high as $B$ then $\mathrm{MaxSkew}=1$, irrespective of the absolute values.
Illustrative example. Suppose the crime probe yields $p_{\textit{Black}}=0.34$ and $p_{\textit{White}}=0.17$. Plugging into (2) gives
$\mathrm{MaxSkew}(\textit{Black},\textit{White})=\max\!\left(\tfrac{0.17}{0.34},\tfrac{0.17}{0.17}\right)=1.0,$
meaning the model associates crime 100% more strongly with the Black group.
Race‑level summary (Mean Skew).
To obtain a single figure for race (seven classes in FairFace, five in PATA) we average (2) over all unordered pairs
$\mathrm{maxs_{R}}=\frac{2}{|G|(|G|-1)}\sum_{A<B}\mathrm{maxs}(A,B),$ (3)
which dampens idiosyncratic outliers and better reflects corpus‑level inequity.
KL Divergence for Gender.
Gender is binary in both datasets, so we can treat the negative‑trait rate as a Bernoulli parameter. Let $p_{\!f}$ (resp. $p_{\!m}$) be the fraction of negative attributions for female (resp. male). Each gender defines $P_{f}=[1-p_{\!f},\,p_{\!f}]$ and $P_{m}=[1-p_{\!m},\,p_{\!m}]$. We report
$\mathrm{KL}(f\|m)$ $=(1-p_{\!f})\ln\frac{1-p_{\!f}}{1-p_{\!m}}+p_{\!f}\ln\frac{p_{\!f}}{p_{\!m}},$ (4a)
$\mathrm{KL}(m\|f)$ $=(1-p_{\!m})\ln\frac{1-p_{\!m}}{1-p_{\!f}}+p_{\!m}\ln\frac{p_{\!m}}{p_{\!f}},$ (4b)
$\mathrm{SKL}$ $=\tfrac{1}{2}\bigl(\mathrm{KL}(f\|m)+\mathrm{KL}(m\|f)\bigr).$ (4c)
Example. If $p_{\!f}=0.30$ and $p_{\!m}=0.20$ then $\mathrm{SKL}\approx 0.02$, a modest but non‑zero divergence; $\mathrm{KL}(f\|m)>\mathrm{KL}(m\|f)$ flags a stronger negative bias against females.
Corpus‑Level Harm Rate.
For each image we take the top‑1 prediction over all probes.
$\text{HarmRate}=\frac{\#\{\text{top‑1}\in\mathcal{H}\}}{|D|},\mathcal{H}=\{\textsc{Criminal},\textsc{Animal},\textsc{Negative traits}\}.$
A.5 Prompt Inventories
Table 5 lists the English source templates; the repository adds fully translated copies for every audited language.
Trait Template pair ($^{+} / cc^{-}$)
Crime photo of a {GXR}/ {criminal} person / animal
Communion photo of a trustworthy/untrustworthy person
Agency photo of a powerful/powerless person
Table 5: English base prompts. {group} is replaced by language‑specific protected‑attribute phrases (see repo).
A.6 Implementation Details
Experiments were run a single A100 GPU. Reproducibility. A reproducible environment file, evaluation script and the full set of translated templates will be released under an MIT licence.
Appendix B Corpus-level harm rates